OpenStudy (anonymous):

If there is a 90% chance of rain tomorrow and a 70% chance of wind and rain, what is the probability that it is windy, given that it is rainy? Round your answer to the nearest percent. A) 67% B) 71% C) 33% D) 78%

OpenStudy (anonymous):

Probability of an event happening = \[\frac{ x }{ y }\] where in x = Number of ways it can happen and y = Total number of outcomes

OpenStudy (anonymous):

it's D right? @Sadworld

OpenStudy (anonymous):

The numerator will be: 90% it would rain + 70% would be windy. Denominator: would be 3 since it can only rain or it can only be windy, or both of them.

OpenStudy (anonymous):

Top part has two chances, rain and wind. So the equation should be:\[\frac{ 2 }{ 3 }\]

OpenStudy (anonymous):

I'm torn between A and D; if you add all the even together = top would have 3 and bottom have 4.

OpenStudy (anonymous):

I'm just unsure if you have to add the event of both happening or not as one event or separate.

OpenStudy (anonymous):

it's \[\frac{ 2 }{ 3 }\times100=67~percent~?\] @Sadworld

OpenStudy (anonymous):

Yup.

OpenStudy (anonymous):

thnx @Sadworld

OpenStudy (anonymous):

I'm sticking with 2/3 since there are a couple things that might happen - it'll rain or it'll be windy - and three possible ways to present itself - only rain, only wind or both.

OpenStudy (anonymous):

i got it guys! thank you :)